Find an equation in slope-intercept form passing through the points (-6.2, 18.1) and (3.4, 18.1)
...any help here would be appreciated!

Answer by Lacey020991(49)

The slope-intercept formula is
y=mx+b
.
First you have to find your slope
The slope, m, of the line is the change in y divided by the change in x.
.
m = (y2-y1) / (x2-x1)
x1= -6.2 ; y1= 18.1
x2= 3.4 ; y2= 18.1
.
m= (18.1-18.1) / (3.4- -6.2)
m= 0 / 9.6
m= 0
.
If your slope would have 9.6 / 0 the slope would have been undefined and it would have been a vertical line, but since it's a slope of zero it's going to be a horizontal line.
.
Now, to find the y-intercept, b, use one of the points to get the x and y values, and solve for b.
.
y1 = mx1 + b
b = y1 - mx1
.
b = 18.1 - 0(-6.2)
b = 18.1
So your equation is :
y = 18.1
.
Your graph will just be a horizontal line at 18.1.